OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 8x = 2y + 5 3x = y + 7 What is the solution set? {(-41/2, -9/2)} {(-9/2, -41/2)} Ø

Vocaloid (vocaloid):

any ideas?

OpenStudy (anonymous):

yes first is correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

find 2y ie 2y = 8x - 5 6x = 2y + 14 6x = 8x - 5 + 14 -2x = 9 x = -9/2 and y = -41/2 is the answer...